OpenStudy (anonymous):

find all the zeros & state the multiplicity y= x^3+x^2-19x+5

OpenStudy (anonymous):

\[(x+5) (x^2-4 x+1)\] so -5 with multiplity 1, then quadratic formual for other two

OpenStudy (anonymous):

actually for \[x^2-4x+1=0\] easier to complete the square \[x^2-4x=-1\] \[(x-2)^2=-1+4=3\] \[x+2=\pm\sqrt{3}\] \[x=-2\pm\sqrt{3}\]

OpenStudy (anonymous):

typo\[x=2\pm\sqrt{3}\]

OpenStudy (anonymous):

what do you mean typo? which part

OpenStudy (anonymous):

dont you see my ''x''= 2+-sqrt3 but satellite73'1 is different than me..

OpenStudy (anonymous):

he made typo..